# Multivariate Probabilistic Time Series Forecasting with Correlated Errors

### v3

@@ 5.1 Evaluation of Predictive Performance
Datasets. We use widely recognized time series benchmarking datasets from GluonTS [21]. The prediction range ($Q$) for each dataset was based on their respective configurations within GluonTS. We performed a sequential split into training, validation, and testing sets for each dataset. Each dataset was standardized using the mean and standard deviation from the training set. Predictions were rescaled to their original values for computing evaluation metrics. Further details on the dataset can be found in Appendix §A.1.
Base probabilistic models. We integrate the proposed method into two distinct autoregressive models: the RNN-based GPVar [2] and the decoder-only Transformer [22]. These models are trained to generate distribution parameters as described in §4. Our approach can be applied to other autoregressive multivariate models with minimal adjustments, provided the final prediction follows a multivariate Gaussian distribution. The implementation is carried out using PyTorch Forecasting [23]. We use lagged time series values and additional features/covariates as input for both models. Details on model training (§A.3), hyperparameter tuning (§A.5), and the base model (§A.6) are provided in Appendix §A.
Dynamic correlation matrix. We introduce a limited number of additional parameters to project the state vector $\mathbf{h}_{t}$ into component weights $w_{m,t}$, which are used to generate the dynamic correlation matrix $\boldsymbol{C}_{t}$. The number of base kernels ($M$) for generating $\boldsymbol{C}_{t}$ and the associated lengthscale set $\{l_{m}\}_{m=1}^{M-1}$ are hyperparameters. We conduct a grid search over $M=2,3,4$ and two sets of lengthscales—$\{0.5,1.5,\dots\}$ and $\{1.0,2.0,\dots\}$. Models with the best validation loss are selected. These distinct lengthscales capture varying correlation decay rates, allowing the model to account for different temporal patterns. The time-varying component weights enable dynamic adaptation to evolving correlation structures over time.
Baselines. We assess the proposed method by comparing it with a baseline model trained without considering error cross-correlation (Eq. (4)). The baseline model is effectively a special case of our model with $\boldsymbol{C}_{t}=\boldsymbol{I}_{D}$. To ensure a straightforward and fair comparison, we align the cross-correlation range ($D$) with the prediction range ($Q$), ensuring identical data sampling processes for both methods. Additionally, we set $P=Q$ following the default configuration in GluonTS. We also provide VAR and GARCH as two naive baseline models (see Appendix §A.4).
Metrics. Continuous Ranked Probability Score (CRPS) [24] is used as the main evaluation metric:
@@
where $F$ is the cumulative distribution function (CDF) of the predicted variable, $z$ is the observation, $Z$ and $Z^{\prime}$ are independent copies of a set of prediction samples associated with the distribution $F$. To evaluate multivariate dependencies in the time series data, we compute $\operatorname{CRPS}_{\text{sum}}$ by first summing both the forecast and ground-truth values across all time series and then calculating the $\operatorname{CRPS}$ over the resulting sums [2, 15, 16]. Other metrics are also available in Appendix §B.1.
Training dynamics. Our approach incurs additional training costs per optimization step due to the more complex likelihood function. As shown in Appendix §B.3, the training time per epoch for models using our method is generally longer than for baseline methods. However, there is no evidence that our method slows down convergence. On the contrary, it accelerates convergence for 5 out of 9 datasets for GPVar.
Benchmark results. The $\operatorname{CRPS}_{\text{sum}}$ results are presented in Table 1. Our method offers an average improvement of 13.79% for GPVar and 6.91% for Transformer. It is important to note that the degree of performance enhancement varies across different base models and datasets, influenced by factors such as the inherent data characteristics and the performance of different model architectures. The alignment between the actual correlation structure and our kernel assumption also plays a pivotal role in our method. Our method also demonstrates consistent improvement across five different metrics, with significant gains in multivariate metrics such as the energy score (Appendix §B.2). Additionally, to provide more insights, we compare the residual autocorrelation and cross-lag correlation with and without applying our method in Appendix §B.4.1, showing that our method effectively reduces cross-correlations in many scenarios. Furthermore, in Appendix §B.4.2, we separate the accuracy improvement over forecast steps for each dataset. The performance improvement is shown to be related to both the absolute time across the temporal span of the dataset (especially for time series with strong periodic patterns) and the relative time over the prediction horizon.
@@ 5.2 Model Interpretation
Our method captures error cross-correlation through the dynamic construction of a covariance matrix. This is achieved by combining kernel matrices with varying lengthscales in a dynamically weighted sum. A small lengthscale corresponds to short-range positive correlations, while a large lengthscale captures positive correlations over longer lags.
@@
In Fig. 3, we depict the dynamic component weights and the resulting autocorrelation function (the first row of the correlation matrix $\boldsymbol{C}_{t}$) for a batch of time series from the $\mathtt{m4\_hourly}$ dataset spanning a four-day window. We also provide the covariance matrix of $\boldsymbol{z}_{t}^{\text{bat}}$ using the correlation matrix and model outputs at a specific time of day. We observe that the component weight $w_{3}$, corresponding to the identity matrix, dominates throughout the entire observation period. This suggests that the error correlation is generally mild over time. This behavior is influenced by the Kronecker structure used to parameterize the covariance over the low-dimensional latent variables $\boldsymbol{r}_{t}$, which assumes that all latent processes share the same autocorrelation structure. Given the Kronecker structure, the model tends to learn the mildest temporal correlation among the time series in a batch.
@@
Moreover, we observe that the dynamic component weights adjust the correlation strengths. Specifically, when the weight assigned to the identity matrix ($w_{3}$) increases, the error process tends to be more independent. In contrast, when the weights assigned to the other kernel matrices ($w_{0}$, $w_{1}$, and $w_{2}$) are larger, the error process becomes more correlated, as the kernel matrices with different lengthscales combine to formulate a specific correlation structure. Fig. 3(a) demonstrates pronounced daily patterns in temporal correlation, particularly when errors exhibit increased correlation around 17:00 each day. The corresponding autocorrelation function is shown in Fig. 3(b). Fig. 3(c) illustrates the corresponding covariance matrix of the associated target variables within the cross-correlation horizon. The diagonal blocks represent the contemporaneous covariance $\boldsymbol{\Sigma}_{t}$ of $\mathbf{z}_{t}$ at each time step, while the off-diagonal blocks represent the cross-covariance $\operatorname{Cov}(\mathbf{z}_{t-\Delta},\mathbf{z}_{t})$ for $\forall\Delta\neq 0$, which can be captured by our approach. The zoomed-in view provides a $3B\times 3B$ region that illustrates the cross-covariance within two lags. We observe that the cross-covariance is most pronounced at lag 1, consistent with the observation in Fig. 3(a) that the component weight $w_{0}$, assigned to the base kernel matrix with lengthscale $l=0.5$, is more pronounced than $w_{1}$ and $w_{2}$.
@@ 6 Conclusion and Broader Impacts
This paper presents a novel approach to addressing error cross-correlation in multivariate probabilistic time series forecasting, specifically for models with autoregressive properties and Gaussian distribution heads. We construct a dynamic covariance matrix using a small set of independent and identically distributed latent temporal processes. These latent processes effectively model temporal correlation and integrate seamlessly into the base model, where the contemporaneous covariance is parameterized by a low-rank-plus-diagonal structure. This approach enables the modeling and prediction of a time-varying covariance matrix for the target time series variable. The experiment results demonstrate its effectiveness in enhancing uncertainty quantification.
Our contributions are two-fold. First, our approach relaxes the time-independent error assumption during the training process for probabilistic forecasting models, addressing the reality that residuals are typically not time-independent. Second, the learned cross-correlation enhances multistep-ahead predictions by refining the distribution output at each forecasting step. These enhancements to existing models have broader implications for fields such as finance, healthcare, and energy, where improved forecasts and uncertainty quantification can lead to better risk management, resource allocation, and operational efficiency.
There are several avenues for future research. Firstly, the Kronecker structure $\boldsymbol{C}_{t}\otimes\mathbf{I}_{R}$ for the covariance matrix of the latent variable $\boldsymbol{r}_{t}^{\text{bat}}$ may be too restrictive for multivariate time series problems. This structure assumes that the rows (i.e., latent temporal processes) in the matrix $\left[\mathbf{r}_{t-D+1},\ldots,\mathbf{r}_{t}\right]$ are identically distributed following $\mathcal{N}\left(\boldsymbol{0},\boldsymbol{C}_{t}\right)$. Exploring more flexible covariance structures, such as employing different $\boldsymbol{C}_{r,t}$ matrices for each of the latent temporal processes, could be a promising direction for further investigation. Secondly, the parameterization of $\boldsymbol{C}_{t}$ could be expanded. Instead of using SE kernels, $\boldsymbol{C}_{t}$ could be parameterized as fully learnable positive definite symmetric Toeplitz matrices. For example, an AR($p$) process has a covariance structure with a Toeplitz form, allowing for the modeling of negative correlations. This alternative approach could offer greater flexibility in capturing complex correlation patterns in multivariate time series data.
@@
We acknowledge the support of the Natural Sciences and Engineering Research Council of Canada (NSERC). V. Z. Zheng acknowledges the support received from the FRQNT B2X Doctoral Scholarship Program.
@@ B Metrics and Additional Results B
B.4 Additional Model Interpretation B.4
B.4.1 Comparison of Residual Correlation B.4.1
B.4.2 Performance Breakdown at Each Forecast Step B.4.2
B.5 Alternative Parametrization of $\boldsymbol{C}_{t}$ B.5
B.5.1 Learnable Lengthscales B.5.1
B.5.2 Using Autocorrelations of an AR$(p)$ process B.5.2
B.6 Qualitative Results on Forecasting B.6
@@ A.1 Datasets
  • $\mathtt{m4\_hourly}$ [26]: Hourly time series data from various domains, covering microeconomics, macroeconomics, finance, industry, demographics, and various other fields, is sourced from the M4-competition.
@@
These datasets are widely employed for benchmarking time series forecasting models, following their default configurations in GluonTS, including granularity, prediction range ($Q$) and the number of rolling evaluations. For each dataset, we performed a sequential split into training, validation, and testing sets, with the temporal length of the validation set matching that of the testing set. The temporal length of the testing set was determined by considering the prediction range and the required number of rolling evaluations. For instance, the testing horizon for the $\mathtt{traffic}$ dataset is computed as $24+7-1=30$ time steps. As a result, the model will predict 24 steps ($Q$) sequentially, with 7 distinct consecutive prediction start timestamps, also known as 7 forecast instances. In our experiments, we align the conditioning range ($P$) with the prediction range ($Q$), maintaining consistency with the default setting in GluonTS. For simplicity, we set the autocorrelation horizon ($D$) to also match the prediction range ($Q$). Essentially, in this paper, we have $P=Q=D$. Each dataset was standardized using the mean and standard deviation from the training set. Predictions were rescaled to their original values for computing evaluation metrics. The statistics of all datasets are summarized in Table 2.
@@ A.2 Multivariate Likelihood with Correlated Errors
The probability density function of a multivariate normal distribution considering autocorrelated errors described in §4.1 is defined in Eq. (14). For simplicity, we omit the subscript $t$ and superscript bat for all notations:
@@
where $B$ is the number of time series in a batch, and the covariance matrix is parameterized as $\boldsymbol{\Sigma}=\boldsymbol{L}(\boldsymbol{C}\otimes\mathbf{I}_{R})\boldsymbol{L}^{\top}+\boldsymbol{E}$, $\boldsymbol{L}\in\mathbb{R}^{DB\times DR}$ the covariance factor, $\boldsymbol{C}\in\mathbb{R}^{D\times D}$ the autocorrelation matrix, $\boldsymbol{E}=\operatorname{diag}(\boldsymbol{d})$, $\boldsymbol{d}\in\mathbb{R}_{+}^{DB}$ the diagonal elements. The bottleneck of evaluating this NLL lies on the calculation of the inverse and determinant of $\boldsymbol{\Sigma}$. Here, we can simplify the calculation using the Sherman–Morrison–Woodbury identity (matrix inversion lemma) and the companion matrix determinant lemma:
@@
where $\boldsymbol{k}=\boldsymbol{L}_{cap}^{-1}\boldsymbol{L}^{\top}\boldsymbol{E}^{-1}\boldsymbol{\eta}$, $\boldsymbol{L}_{cap}$ is the Cholesky factor of the capacitance matrix $\left(\left(\boldsymbol{C}\otimes\mathbf{I}_{R}\right)^{-1}+\boldsymbol{L}^{\top}\boldsymbol{E}^{-1}\boldsymbol{L}\right)$, the computing of $\boldsymbol{k}=\boldsymbol{L}_{cap}^{-1}\boldsymbol{L}^{\top}\boldsymbol{E}^{-1}\boldsymbol{\eta}$ can be fastly resolved by solving a linear system of equations $\boldsymbol{L}_{cap}\boldsymbol{k}=\boldsymbol{L}^{\top}\boldsymbol{E}^{-1}\boldsymbol{\eta}$. Since $\boldsymbol{E}$ is a diagonal matrix, the only matrix inverse we need to calculate in Eq. (17) is $\left(\boldsymbol{C}\otimes\mathbf{I}_{R}\right)^{-1}$, which can be further simplified as $\boldsymbol{C}^{-1}\otimes\mathbf{I}_{R}$. Recall $\boldsymbol{C}$ is a $D\times D$ autocorrelation matrix. Therefore, it is much easier than calculating the inverse of $\boldsymbol{\Sigma}$ which is a $DB\times DB$ matrix, and the computational cost does not scale with the number of time series $B$ in a batch.

### v4

@@ 5.1 Evaluation of Predictive Performance
Datasets. We use widely recognized time series benchmarking datasets from GluonTS [24]. The prediction range ($Q$) for each dataset follows the configurations provided by GluonTS. We applied a sequential split into training, validation, and testing sets for each dataset. Each dataset was standardized using the mean and standard deviation from the training set, and predictions were rescaled to their original values for evaluation. Further details on the datasets can be found in Appendix §A.1.
Base probabilistic models. We integrated the proposed method into two distinct autoregressive models: the RNN-based GPVar [3] and the decoder-only Transformer [25]. These models are trained to generate distribution parameters as described in §4. Our approach can be applied to other autoregressive multivariate models with minimal adjustments, provided the final prediction follows a multivariate Gaussian distribution. The implementation is based on using PyTorch Forecasting [26]. Both models use lagged time series values and additional features or covariates as inputs. Details on model training (§A.3), hyperparameter tuning (§A.5), and the base model (§A.6) are provided in Appendix §A. The code is available at https://github.com/rottenivy/mv_pts_correlatederr.
Dynamic correlation matrix. We introduce a limited number of additional parameters to project the state vector $\mathbf{h}_{t}$ into component weights $w_{m,t}$, which are used to generate the dynamic correlation matrix $\boldsymbol{C}_{t}$. The number of base kernels ($M$) for generating $\boldsymbol{C}_{t}$ and the associated lengthscale set $\{l_{m}\}_{m=1}^{M-1}$ are treated as hyperparameters. We perform a grid search over $M=2,3,4$ and two sets of lengthscales—$\{0.5,1.5,\dots\}$ and $\{1.0,2.0,\dots\}$. Models with the best validation loss are selected. These different lengthscales capture varying correlation decay rates, enabling the model to account for different temporal patterns. The time-varying component weights enable dynamic adaptation to changing correlation structures over time.
Baselines. We evaluate the proposed method by comparing it with a baseline model trained without accounting for error cross-correlation (Eq. (4)). The baseline model represents a special case of our model with $\boldsymbol{C}_{t}=\boldsymbol{I}_{D}$. To ensure a straightforward and fair comparison, we align the cross-correlation range ($D$) with the prediction range ($Q$), ensuring identical data sampling processes for both methods. Additionally, we set $P=Q$ following the default configuration in GluonTS. We also include VAR and GARCH as naive baseline models (see Appendix §A.4).
Metrics. We use the Continuous Ranked Probability Score (CRPS) [27] as the main metric:
@@
where $F$ is the cumulative distribution function (CDF) of the predicted variable, $z$ is the observation, $Z$ and $Z^{\prime}$ are independent copies of the prediction samples associated with the distribution $F$. To evaluate multivariate dependencies in the time series data, we compute $\operatorname{CRPS}_{\text{sum}}$ by first summing both the forecast and ground-truth values across all time series and then calculating the $\operatorname{CRPS}$ over the resulting sums [3, 16, 17]. As $\operatorname{CRPS}_{\text{sum}}$ may overlook model performance on individual dimensions [28], we also report additional metrics, e.g., the energy score [27, 29], in Appendix §B.1.
Training dynamics. Our approach incurs additional training costs per optimization step due to the more complex likelihood function. As shown in Appendix §B.3, the training time per epoch for models using our method is generally longer than that of baseline methods. However, our parameterization allows for scalability to large time series datasets by using a small random subset of time series at each optimization step during training.
@@
Benchmark results. The $\operatorname{CRPS}_{\text{sum}}$ results are presented in Table 1. Our method achieves an average improvement of 13.79% for GPVar and 6.91% for the Transformer model. It is important to note that the degree of performance enhancement varies across different base models and datasets, influenced by factors such as the inherent data characteristics and the performance of different model architectures. The alignment between the actual correlation structure and our kernel assumption also plays a crucial role in the effectiveness of our method. Additionally, our approach demonstrates consistent improvements across five different metrics, with significant gains in multivariate metrics such as the energy score (Appendix §B.2).
To provide further insights, we compare the residual autocorrelation and cross-lag correlation with and without applying our method in Appendix §B.5.1, showing that our method effectively reduces cross-correlations in many scenarios. We use ACF plot comparisons to illustrate the reduction in autocorrelation and cross-correlation plot comparisons to demonstrate the decrease in cross-lag correlation. The residuals generated by the model with our method exhibit weaker cross-correlations, which is particularly enhanced by the calibration process during prediction (§4.2).
Furthermore, Appendix §B.5.2 separates the accuracy improvement over forecast steps for each dataset. The performance improvement is shown to be related to both the absolute time across the temporal span of the dataset (especially for time series with strong periodic patterns) and the relative time over the prediction horizon.
5.2 Model Interpretation
@@
Our method captures error cross-correlation through the dynamic construction of a covariance matrix, achieved by combining kernel matrices with varying lengthscales in a dynamically weighted sum. A small lengthscale corresponds to short-range positive correlations, while a large lengthscale captures positive correlations over longer lags.
@@
In Fig. 4, we depict the dynamic component weights and the resulting autocorrelation function (the first row of the correlation matrix $\boldsymbol{C}_{t}$) for a batch of time series from the $\mathtt{m4\_hourly}$ dataset spanning a four-day window. We also provide the covariance matrix of $\boldsymbol{z}_{t}^{\text{bat}}$ using the correlation matrix and model outputs at a specific time of day. The component weight $w_{3}$, corresponding to the identity matrix, dominates throughout the observation period. This suggests that the error correlation is generally mild over time. This behavior is influenced by the Kronecker structure used to parameterize the covariance over the low-dimensional latent variables $\boldsymbol{r}_{t}$, which assumes all latent processes share the same autocorrelation structure. Given the Kronecker structure, the model tends to learn the mildest temporal correlation among the time series in a batch.
@@
Moreover, we observe that the dynamic component weights adjust the correlation strengths. Specifically, when the weight assigned to the identity matrix ($w_{3}$) increases, the error process tends to be more independent. In contrast, when the weights assigned to the other kernel matrices ($w_{0}$, $w_{1}$, and $w_{2}$) are larger, the error process becomes more correlated, as the kernel matrices with different lengthscales combine to formulate a specific correlation structure. Fig. 4(a) demonstrates pronounced daily patterns in temporal correlation, particularly when errors exhibit increased correlation around 17:00 each day. The corresponding autocorrelation function is shown in Fig. 4(b). Fig. 4(c) illustrates the corresponding covariance matrix of the associated target variables within the cross-correlation horizon. The diagonal blocks represent the contemporaneous covariance $\boldsymbol{\Sigma}_{t}$ of $\mathbf{z}_{t}$ at each time step, while the off-diagonal blocks capture the cross-covariance $\operatorname{Cov}(\mathbf{z}_{t-\Delta},\mathbf{z}_{t})$ for $\forall\Delta\neq 0$, effectively modeled by our approach. The zoomed-in view provides a $3B\times 3B$ region that illustrates the cross-covariance within two lags. We observe that the cross-covariance is most pronounced at lag 1, consistent with the observation in Fig. 4(a) that the component weight $w_{0}$, assigned to the base kernel matrix with lengthscale $l=0.5$, is more pronounced than $w_{1}$ and $w_{2}$.
6 Discussion
In this section, we discuss factors that influence the performance of our method. Specifically, we highlight the effectiveness of our model in long-term forecasting across various scenarios. We also discuss the effect of scaling up to larger batch sizes during prediction. Additionally, we examine the impact of non-Gaussian errors on model performance.
Long-term forecasting. The advantage of modeling error correlation can vary in long-term forecasting, especially in autoregressive predictions where errors accumulate and propagate over time. Using residuals from previous time steps to calibrate forecasts may be beneficial for non-stationary segments of the time series. However, for time series with strong periodic effects, the model may also rely on seasonal lags. As shown in Fig. 21 and Fig. 22 of the Appendix, the advantage of modeling error correlation can decrease in longer-term forecasts compared to shorter-term forecasts for some datasets with strong periodic effects (e.g., the $\mathtt{traffic}$ dataset in Fig. 21). It is not necessarily true that the advantage diminishes for long-horizon predictions, as the effectiveness of our method depends on the quality of predictions during inference. In cases where the model provides accurate long-term forecasts, the benefit of modeling correlated errors may be less pronounced.
Scalability. Increasing the number of time series $B$ in a batch leads to higher training costs. Because the model requires numerous iterations over the dataset for optimization, using a large $B$ during training is not feasible. However, during prediction, the batch size can be increased to leverage more information. This may enhance both prediction accuracy and error calibration, provided sufficient memory is available. We demonstrate the effect of increasing batch size during prediction in Appendix §B.4 through additional experiments. Both models, with and without our method, show improvement from increased batch sizes during prediction, as reflected by a decrease in $\operatorname{CRPS}_{\text{sum}}$.
Non-Gaussian errors. For the baseline model, assuming Gaussian errors may lead to model misspecification, resulting in more correlated residuals. To address this issue, we also trained the baseline models using the likelihood of a multivariate $t$-distribution; the results are presented in Table 15 of the Appendix. Although using an alternative distribution can lead to better performance on some datasets without our method, we observed that our method effectively closes the performance gap when the $t$-distribution outperforms the Gaussian assumption. We chose the Gaussian distribution for its beneficial properties, including its marginalization rule and well-defined conditional distributions, both essential for statistically consistent model training and reliable inference. Thus, a more effective approach could involve first transforming the original observations into Gaussian-distributed data using a Gaussian Copula [3], followed by applying our method.
@@ 7 Conclusion and Broader Impacts
This paper presents a novel approach for addressing error cross-correlation in multivariate probabilistic time series forecasting, specifically for models with autoregressive properties and Gaussian distribution outputs. We construct a dynamic covariance matrix using a small set of independent and identically distributed latent temporal processes. These latent processes effectively model temporal correlation and integrate seamlessly into the base model, where the contemporaneous covariance is parameterized by a low-rank-plus-diagonal structure. This approach enables the modeling and prediction of a time-varying covariance matrix for the target time series variables. The experimental results demonstrate its effectiveness in enhancing uncertainty quantification.
Our contributions are two-fold. First, our approach relaxes the time-independent error assumption during the training process for probabilistic forecasting models, addressing the reality that residuals are typically time-dependent. Second, the learned cross-correlation improves multistep-ahead predictions by refining the distribution output at each forecasting step. These enhancements to existing models have broader implications for fields such as finance, healthcare, and energy, where improved forecasts and uncertainty quantification can lead to more informed decisions.
There are several avenues for future research. First, the Kronecker structure $\boldsymbol{C}_{t}\otimes\mathbf{I}_{R}$ for the covariance matrix of the latent variable $\boldsymbol{r}_{t}^{\text{bat}}$ may be too restrictive for multivariate time series problems. Exploring more flexible covariance structures, such as employing different $\boldsymbol{C}_{r,t}$ matrices for each latent temporal process as in the linear model of coregionalization (LMC, [22, 23]), could be a promising direction for further investigation. Second, the parameterization of $\boldsymbol{C}_{t}$ could be expanded. Instead of using SE kernels, $\boldsymbol{C}_{t}$ could be parameterized as fully learnable positive definite symmetric Toeplitz matrices. For example, an AR($p$) process has a covariance structure in Toeplitz form, allowing for the modeling of negative correlations. This alternative approach could offer greater flexibility in capturing complex correlation patterns in multivariate time series data.
@@
We acknowledge the support from the Natural Sciences and Engineering Research Council (NSERC) of Canada (Discovery Grant). Vincent Zhihao Zheng also acknowledges the support received from the FRQNT B2X Doctoral Scholarship Program.
@@ B Metrics and Additional Results B
B.4 Effect of the Number of Time Series during Prediction B.4
B.5 Additional Model Interpretation B.5
B.5.1 Comparison of Residual Correlation B.5.1
B.5.2 Performance Breakdown at Each Forecast Step B.5.2
B.6 Alternative Parametrization of $\boldsymbol{C}_{t}$ B.6
B.6.1 Learnable Lengthscales B.6.1
B.6.2 Using Autocorrelations of an AR$(p)$ process B.6.2
B.7 Alternative Error Assumptions B.7
B.8 Qualitative Results on Forecasting B.8
@@ A.1 Datasets
  • $\mathtt{m4\_hourly}$ [31]: Hourly time series data from various domains, covering microeconomics, macroeconomics, finance, industry, demographics, and various other fields, are sourced from the M4-competition.
@@
These datasets are widely employed for benchmarking time series forecasting models, following their default configurations in GluonTS, including granularity, prediction range ($Q$), and the number of rolling evaluations. For each dataset, we performed a sequential split into training, validation, and testing sets, with the temporal length of the validation set matching that of the testing set. The temporal length of the testing set was determined by considering the prediction range and the required number of rolling evaluations. For instance, the testing horizon for the $\mathtt{traffic}$ dataset is computed as $24+7-1=30$ time steps. As a result, the model will predict 24 steps ($Q$) sequentially, with 7 distinct consecutive prediction start timestamps, also known as 7 forecast instances. In our experiments, we align the conditioning range ($P$) with the prediction range ($Q$), maintaining consistency with the default setting in GluonTS. For simplicity, we set the autocorrelation horizon ($D$) to also match the prediction range ($Q$). Essentially, in this paper, we have $P=Q=D$. Each dataset was standardized using the mean and standard deviation from the training set. Predictions were rescaled to their original values for computing evaluation metrics. The statistics of all datasets are summarized in Table 2.
@@ A.2 Multivariate Likelihood with Correlated Errors
The probability density function of a multivariate normal distribution with autocorrelated errors, as described in §4.1, is defined in Eq. (14). For simplicity, we omit the subscript $t$ and superscript bat for all notations:
@@
where $B$ is the number of time series in a batch. The covariance matrix is parameterized as $\boldsymbol{\Sigma}=\boldsymbol{L}(\boldsymbol{C}\otimes\mathbf{I}_{R})\boldsymbol{L}^{\top}+\boldsymbol{E}$. In this parameterization, $\boldsymbol{L}\in\mathbb{R}^{DB\times DR}$ is the covariance factor, $\boldsymbol{C}\in\mathbb{R}^{D\times D}$ is the autocorrelation matrix, $\boldsymbol{E}=\operatorname{diag}(\boldsymbol{d})$, and $\boldsymbol{d}\in\mathbb{R}_{+}^{DB}$ are the diagonal elements. The bottleneck in evaluating this NLL lies in the calculation of the inverse and determinant of $\boldsymbol{\Sigma}$. Therefore, we can simplify the calculation using the Sherman–Morrison–Woodbury identity (matrix inversion lemma) and the companion matrix determinant lemma:
@@
where $\boldsymbol{k}=\boldsymbol{L}_{cap}^{-1}\boldsymbol{L}^{\top}\boldsymbol{E}^{-1}\boldsymbol{\eta}$. $\boldsymbol{L}_{cap}$ is the Cholesky factor of the capacitance matrix $\left(\left(\boldsymbol{C}\otimes\mathbf{I}_{R}\right)^{-1}+\boldsymbol{L}^{\top}\boldsymbol{E}^{-1}\boldsymbol{L}\right)$. The computation of $\boldsymbol{k}$ can be efficiently resolved by solving the linear system of equations $\boldsymbol{L}_{cap}\boldsymbol{k}=\boldsymbol{L}^{\top}\boldsymbol{E}^{-1}\boldsymbol{\eta}$. Since $\boldsymbol{E}$ is a diagonal matrix, the only matrix inverse we need to calculate in Eq. (17) is $\left(\boldsymbol{C}\otimes\mathbf{I}_{R}\right)^{-1}$, which can be further simplified as $\boldsymbol{C}^{-1}\otimes\mathbf{I}_{R}$. Recall that $\boldsymbol{C}$ is a $D\times D$ autocorrelation matrix. Therefore, calculating its inverse is much easier than computing the inverse of $\boldsymbol{\Sigma}$, which is a $DB\times DB$ matrix. Moreover, the computational cost does not scale with the number of time series $B$ in a batch.